# Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies

### v2

@@ Acknowledgments
Appendix A Additional Information
In the literature, positivity violations are typically classified into two types—structural and practical—usually in the context of binary interventions [38, 27, 37]. Analogous definitions can be formulated for continuous interventions.
A.1 Near positivity violations
Near violations occur when the positivity assumption is almost, but not completely, violated. Here, the probability of receiving certain treatment levels given specific confounder combinations is very close to zero but not exactly zero. Although not a strict violation, the consequences are similar: bias and loss of precision in causal estimation [24]. Estimators become unstable in these regions, since causal methods are particularly sensitive to sparse data [24].
A.2 Structural positivity violations
Structural violations arise when, due to the design or data structure, some subgroups of individuals (defined by their confounders) have no chance of receiving certain treatment values. Formally, $f_{A|\mathbf{L}}(a|\mathbf{L})=0$ for these combinations.
A.3 Practical positivity violations
Also called stochastic violations, these occur when a treatment is clinically possible for some confounder strata but is unobserved in the data, leading to $\hat{f}_{A|\mathbf{L}}(a|\mathbf{L})=0$. For continuous interventions this notion is more subtle: exact intervention values are never observed, so whether $\hat{f}(a|\mathbf{L})=0$ depends on how the conditional density is estimated.
A.4 Theoretical implications for plug-in estimation
Positivity violations—whether structural (deterministic absence of support) or practical (sparse data)—create fundamental difficulties for plug-in estimation of the causal dose–response curve (CDRC) in Equation 6. To formalize these issues, let
$Q_{0}(a,\mathbf{l}):=E_{P_{0}}(Y\mid A=a,\mathbf{L}=\mathbf{l})$
denote the true conditional outcome regression. Under standard identification assumptions, the CDRC is identified by the g-formula:
@@ A.4 Theoretical implications for plug-in estimation
Step 1: Estimation of the regression function.
In practice, one estimates $Q_{0}(a,\mathbf{l})=E_{P_{0}}(Y\mid A=a,\mathbf{L}=\mathbf{l})$ by fitting a regression function $\hat{Q}(A,\mathbf{L})$, for example via least squares. The corresponding population risk of a candidate function $Q$ is
@@
By the variance decomposition identity,
$R(Q)=E_{P_{0}}\!\left[(Q_{0}(A,\mathbf{L})-Q(A,\mathbf{L}))^{2}\right]+E_{P_{0}}\!\left[\mathrm{Var}(Y\mid A,\mathbf{L})\right].$
Since the second term does not depend on $Q$, minimizing (22) is equivalent to minimizing
$\int\big(Q_{0}(A,\mathbf{L})-Q(A,\mathbf{L})\big)^{2}\,dP_{A,\mathbf{L}}.$ (23)
Thus, estimation implicitly weights squared errors by the observed joint density $f_{A,\mathbf{L}}(a,\mathbf{l})=f_{A|\mathbf{L}}(a\mid\mathbf{l})f_{\mathbf{L}}(\mathbf{l})$. Equivalently, the contribution of a given stratum $\mathbf{l}$ to the risk is scaled by its marginal probability $f_{\mathbf{L}}(\mathbf{l})$, and within each stratum the influence of treatment level $a$ is weighted by its conditional density $f_{A|\mathbf{L}}(a\mid\mathbf{l})$. Hence, regions with low support (small $f_{A|\mathbf{L}}(a\mid\mathbf{l})$) contribute negligibly to the risk, and therefore exert little influence on the fitted regression $\hat{Q}$.
@@
The plug-in estimator of (21) is
$\hat{m}(a)=E_{P_{n}}\!\left[\hat{Q}(a,\mathbf{L})\right],$
where $P_{n}$ is the empirical distribution of $\mathbf{L}$ and $\hat{Q}$ is the fitted regression. Its mean squared error decomposes as
$\text{MSE}(\hat{m}(a))=E_{P_{0}}\!\left[\big(\hat{Q}(a,\mathbf{L})-Q_{0}(a,\mathbf{L})\big)^{2}\right]=\int\big(\hat{Q}(a,\mathbf{l})-Q_{0}(a,\mathbf{l})\big)^{2}\,dP_{\mathbf{L}}(\mathbf{l}).$ (24)
The mismatch:
Equations (22) and (24) highlight a fundamental mismatch:
  • The regression step minimizes loss with respect to $P_{A,\mathbf{L}}$, effectively weighting errors by $f_{A|\mathbf{L}}(a\mid\mathbf{l})f_{\mathbf{L}}(\mathbf{l})$.
  • The target parameter averages errors with respect to the marginal distribution $P_{\mathbf{L}}$, so that each stratum contributes in proportion to its prevalence in the population, regardless of whether the treatment level $A=a$ is well supported in that stratum.
Formally, the regression risk is
$R(g)=\int\big(Q_{0}(a,\mathbf{l})-g(a,\mathbf{l})\big)^{2}f_{A|\mathbf{L}}(a\mid\mathbf{l})\,f_{\mathbf{L}}(\mathbf{l})\,da\,d\mathbf{l},$
whereas the CDRC error criterion is
$\text{MSE}(\hat{m}(a))=\int\big(Q_{0}(a,\mathbf{l})-\hat{Q}(a,\mathbf{l})\big)^{2}f_{\mathbf{L}}(\mathbf{l})\,d\mathbf{l}.$
Thus, if $f_{A|\mathbf{L}}(a\mid\mathbf{l})\approx 0$, errors in approximating $Q_{0}(a,\mathbf{l})$ are essentially ignored when fitting $\hat{Q}$, but those same errors can dominate the plug-in bias in (24) because they are weighted by $f_{\mathbf{L}}(\mathbf{l})$ rather than $f_{A|\mathbf{L}}(a\mid\mathbf{l})$.
Implication. This discrepancy explains why plug-in g-computation is highly sensitive to positivity violations. Even if the regression is consistent in regions with support, small errors in sparse or unsupported strata—where extrapolation is unavoidable—are not controlled by the empirical risk but nevertheless contribute directly to bias in $\hat{m}(a)$.
@@ B.3 Data Generating Process for Simulation 3
Appendix C Estimate the Conditional Density via Hazard Binning
Consider discretizing the observed variable $a$ into $T$ bins using an appropriate binning procedure. Given this, we can estimate the probability of $a$ falling into the interval $\alpha_{t-1},\alpha_{t}),\ t=1,\cdots,T$, using a standard classification model. Within each bin $\alpha_{t-1},\alpha_{t})$, the density is assumed to undergo a uniform transformation, leading to the following expressions:
$P(a\in[\alpha_{t-1},\alpha_{t})|\mathbf{L})$ $=P(a\in[\alpha_{t-1},\alpha_{t})|A\geq\alpha_{t-1},\mathbf{L})\times$

### v3

@@ Acknowledgments
Appendix A Additional Information on Positivity Violation
In the literature, positivity violations are typically classified into structural and practical forms, usually for binary interventions [39, 27, 38]. The same ideas extend to continuous interventions.
Near positivity violations.
Near violations occur when positivity is not exactly violated but is close to failing: some treatment levels have conditional probability (or density) very close to zero for certain confounder patterns. Although not a strict violation, the practical consequences are similar—loss of precision and potential instability in causal estimation, especially in sparse regions [24].
Structural positivity violations.
Structural violations arise when some treatment values are impossible for certain confounder strata by design or mechanism. In continuous-treatment notation, this corresponds to
$f_{A\mid\mathbf{L}}(a\mid\mathbf{L})=0$
for those strata and treatment values.
Practical positivity violations.
Practical (stochastic) violations occur when treatment values are clinically possible but poorly represented in the observed sample. In finite samples this is reflected through estimated support, e.g. $\hat{f}_{A\mid\mathbf{L}}(a\mid\mathbf{L})$ near zero. For continuous treatments, this depends on density-estimation choices because exact treatment values are not repeatedly observed.
Theoretical implications for plug-in estimation.
Positivity violations—whether structural (deterministic lack of support) or practical (data sparsity)—create core challenges for plug-in estimation of the causal dose–response curve (CDRC) in Equation (6). Define
$Q_{0}(a,\mathbf{l}):=E_{P_{0}}(Y\mid A=a,\mathbf{L}=\mathbf{l}).$
Under standard identification assumptions, the CDRC is
@@ Appendix A Additional Information on Positivity Violation
Step 1: Outcome-regression fitting.
In practice, $Q_{0}$ is estimated by fitting $\hat{Q}(A,\mathbf{L})$ (e.g., least squares). For a candidate $Q$, the population risk is
@@
Using variance decomposition:
$R(Q)=E_{P_{0}}\!\left[(Q_{0}(A,\mathbf{L})-Q(A,\mathbf{L}))^{2}\right]+E_{P_{0}}\!\left[\operatorname{Var}(Y\mid A,\mathbf{L})\right].$
Hence minimizing (17) is equivalent to minimizing
$\int\big(Q_{0}(a,\mathbf{l})-Q(a,\mathbf{l})\big)^{2}\,dP_{A,\mathbf{L}}(a,\mathbf{l}).$
So squared errors are weighted by
$f_{A,\mathbf{L}}(a,\mathbf{l})=f_{A\mid\mathbf{L}}(a\mid\mathbf{l})\,f_{\mathbf{L}}(\mathbf{l}).$
Therefore, low-support regions (small $f_{A\mid\mathbf{L}}(a\mid\mathbf{l})$) have little influence on the fitted regression.
@@
The plug-in estimator is
$\hat{m}(a)=E_{P_{n}}\!\left[\hat{Q}(a,\mathbf{L})\right].$
Its mean-squared error at fixed $a$ is
$\mathrm{MSE}\!\big(\hat{m}(a)\big)=E_{P_{0}}\!\left[\big(\hat{Q}(a,\mathbf{L})-Q_{0}(a,\mathbf{L})\big)^{2}\right]=\int\big(\hat{Q}(a,\mathbf{l})-Q_{0}(a,\mathbf{l})\big)^{2}\,dP_{\mathbf{L}}(\mathbf{l}).$ (18)
Key mismatch.
Equations (17) and (18) imply:
  • Regression fitting minimizes error under $P_{A,\mathbf{L}}$, i.e., weighted by $f_{A\mid\mathbf{L}}(a\mid\mathbf{l})\,f_{\mathbf{L}}(\mathbf{l})$.
  • The CDRC target at fixed $a$ averages over $P_{\mathbf{L}}$, i.e., weighted only by $f_{\mathbf{L}}(\mathbf{l})$, regardless of local support at $a$.
Equivalently,
$R(Q)=\int\big(Q_{0}(a,\mathbf{l})-Q(a,\mathbf{l})\big)^{2}f_{A\mid\mathbf{L}}(a\mid\mathbf{l})f_{\mathbf{L}}(\mathbf{l})\,da\,d\mathbf{l},$
whereas
$\mathrm{MSE}\!\big(\hat{m}(a)\big)=\int\big(Q_{0}(a,\mathbf{l})-\hat{Q}(a,\mathbf{l})\big)^{2}f_{\mathbf{L}}(\mathbf{l})\,d\mathbf{l}.$
Thus, when $f_{A\mid\mathbf{L}}(a\mid\mathbf{l})\approx 0$, errors in $Q_{0}(a,\mathbf{l})$ receive little penalty during regression fitting but can still contribute substantially to plug-in bias in (18).
Implication. This mismatch explains why plug-in g-computation can be sensitive to positivity violations: extrapolation errors in sparse or unsupported strata may be weakly controlled by regression risk yet still strongly affect estimation of $m(a)$.
@@ B.3 Data Generating Process for Simulation 3
Appendix C Computing the HDR Cutoff via a Quantile of the Density-Value Distribution
C.1 Highest density cutoff and Highest density region
Fix $\alpha\in(0,1)$ and $\mathbf{l}$. Let $f_{0}(a\mid\mathbf{l})$ denote the conditional density of $A\mid\mathbf{L}=\mathbf{l}$ on a compact support $\mathcal{A}\subset\mathbb{R}$. Define the induced density-value random variable [20]
$Z_{\mathbf{l}}=f_{0}(A\mid\mathbf{l}),\qquad A\sim P_{A\mid\mathbf{L}=\mathbf{l}}.$
For any threshold $t\in\mathbb{R}$, the event $\{Z_{\mathbf{l}}\geq t\}$ is equivalent to $\{f_{0}(A\mid\mathbf{l})\geq t\}$, and therefore
$P\!\big(Z_{\mathbf{l}}\geq t\mid\mathbf{L}=\mathbf{l}\big)=P\!\big(f_{0}(A\mid\mathbf{l})\geq t\mid\mathbf{L}=\mathbf{l}\big)=P\!\big(A\in\{a\in\mathcal{A}:\ f_{0}(a\mid\mathbf{l})\geq t\}\mid\mathbf{L}=\mathbf{l}\big).$
Thus, selecting $t$ as a quantile of $Z_{\mathbf{l}}$ yields a density threshold whose associated upper level set has a prescribed conditional probability content. We define the cutoff as the $(1-\alpha)$-quantile of $Z_{\mathbf{l}}$:
$f_{\alpha}(\mathbf{l})=\inf\Big\{t\in\mathbb{R}:\ P\!\big(Z_{\mathbf{l}}\leq t\mid\mathbf{L}=\mathbf{l}\big)\geq 1-\alpha\Big\}.$ (19)
The highest density region (HDR) at level $\alpha$ is then given by the corresponding upper level set
$\mathcal{A}_{\alpha}(\mathbf{l};P_{0})=\big\{a\in\mathcal{A}:\ f_{0}(a\mid\mathbf{l})\geq f_{\alpha}(\mathbf{l})\big\}.$ (20)
By construction,
$P\!\left(A\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0})\mid\mathbf{L}=\mathbf{l}\right)=P\!\left(Z_{\mathbf{l}}\geq f_{\alpha}(\mathbf{l})\mid\mathbf{L}=\mathbf{l}\right)\;\geq\;\alpha,$
with equality whenever there is no probability mass at the cutoff, i.e., $P\!\left(Z_{\mathbf{l}}=f_{\alpha}(\mathbf{l})\mid\mathbf{L}=\mathbf{l}\right)=0$. In the presence of ties ($P(Z_{\mathbf{l}}=f_{\alpha}(\mathbf{l})\mid\mathbf{L}=\mathbf{l})>0$), the closed-set convention in (20) yields probability content at least $\alpha$, which is the natural analogue of quantile ties in the continuous/discrete boundary case.
C.2 Grid-based sample analogue
Let $\hat{f}_{n}(a\mid\mathbf{l})$ be an estimated conditional density. Let $\overline{\mathbf{a}}^{\,\mathrm{hdr}}=(a^{\mathrm{hdr}}_{1},\dots,a^{\mathrm{hdr}}_{m_{\mathrm{hdr}}})^{\top}$ be a strictly increasing grid on $\mathcal{A}$, i.e., $a^{\mathrm{hdr}}_{1}<\cdots<a^{\mathrm{hdr}}_{m_{\mathrm{hdr}}}$. Define grid cell boundaries $(b_{k})_{k=1}^{m_{\mathrm{hdr}}+1}$ by
$b_{1}:=a^{\mathrm{hdr}}_{1},\qquad b_{k}:=\frac{a^{\mathrm{hdr}}_{k-1}+a^{\mathrm{hdr}}_{k}}{2}\ \ (k=2,\dots,m_{\mathrm{hdr}}),\qquad b_{m_{\mathrm{hdr}}+1}:=a^{\mathrm{hdr}}_{m_{\mathrm{hdr}}}.$
Let the associated partition of $a^{\mathrm{hdr}}_{1},a^{\mathrm{hdr}}_{m_{\mathrm{hdr}}}]$ be given by
$I_{k}:=[b_{k},b_{k+1})\quad(k=1,\dots,m_{\mathrm{hdr}}-1),\qquad I_{m_{\mathrm{hdr}}}:=[b_{m_{\mathrm{hdr}}},b_{m_{\mathrm{hdr}}+1}],$
and define the cell widths as Lebesgue lengths
$w_{k}:=\lambda(I_{k})=b_{k+1}-b_{k},\qquad k=1,\dots,m_{\mathrm{hdr}}.$
On this grid, define the normalized cell masses
$\hat{p}_{n,k}(\mathbf{l})=\frac{\hat{f}_{n}(a^{\mathrm{hdr}}_{k}\mid\mathbf{l})\,w_{k}}{\sum_{r=1}^{m_{\mathrm{hdr}}}\hat{f}_{n}(a^{\mathrm{hdr}}_{r}\mid\mathbf{l})\,w_{r}},\qquad k=1,\dots,m_{\mathrm{hdr}}.$ (21)
Equivalently, $\{\hat{p}_{n,k}(\mathbf{l})\}_{k=1}^{m_{\mathrm{hdr}}}$ defines a discrete distribution on the HDR grid $\{a^{\mathrm{hdr}}_{1},\ldots,a^{\mathrm{hdr}}_{m_{\mathrm{hdr}}}\}$ (or, equivalently, on the partition $\{I_{k}\}_{k=1}^{m_{\mathrm{hdr}}}$). Let $\hat{A}_{\mathbf{l}}$ denote a discrete random variable such that
$P\!\left(\hat{A}_{\mathbf{l}}=a^{\mathrm{hdr}}_{k}\mid\mathbf{L}=\mathbf{l}\right)=\hat{p}_{n,k}(\mathbf{l}),\qquad k=1,\ldots,m_{\mathrm{hdr}}.$
Define the corresponding estimated density-value random variable by
$\hat{Z}_{\mathbf{l}}:=\hat{f}_{n}\!\left(\hat{A}_{\mathbf{l}}\mid\mathbf{l}\right).$
Its conditional distribution function is
$P\!\left(\hat{Z}_{\mathbf{l}}\leq t\mid\mathbf{L}=\mathbf{l}\right)=\sum_{k=1}^{m_{\mathrm{hdr}}}\hat{p}_{n,k}(\mathbf{l})\,\mathbbm{1}\!\left\{\hat{f}_{n}(a^{\mathrm{hdr}}_{k}\mid\mathbf{l})\leq t\right\},\qquad t\in\mathbb{R}.$
Then $\hat{f}_{n,\alpha}(\mathbf{l})$ is the $(1-\alpha)$-quantile of $\hat{Z}_{\mathbf{l}}$:
$\hat{f}_{n,\alpha}(\mathbf{l})=\inf\Big\{t\in\mathbb{R}:\ \sum_{k=1}^{m_{\mathrm{hdr}}}\hat{p}_{n,k}(\mathbf{l})\,\mathbbm{1}\!\big\{\hat{f}_{n}(a^{\mathrm{hdr}}_{k}\mid\mathbf{l})\leq t\big\}\geq 1-\alpha\Big\}.$ (22)
The estimated supported set on the HDR grid is
$\hat{\mathcal{A}}^{\,\mathrm{hdr}}_{n,\alpha}(\mathbf{l})=\Big\{a^{\mathrm{hdr}}_{k}\in\mathcal{A}:\ \hat{f}_{n}(a^{\mathrm{hdr}}_{k}\mid\mathbf{l})\geq\hat{f}_{n,\alpha}(\mathbf{l})\Big\}.$ (23)
Remark C.1 (Quality of the grid approximation).
The quality of the HDR approximation depends on both density-estimation error and grid discretization error.
Let
$\Delta_{n}(\mathbf{l}):=\sup_{a\in\mathcal{A}}\big|\hat{f}_{n}(a\mid\mathbf{l})-f_{0}(a\mid\mathbf{l})\big|,$
and let $\delta_{m}:=\max_{1\leq k\leq m_{\mathrm{hdr}}}w_{k}$ denote the maximum cell width of the HDR-grid partition. Under compact support and regularity of $f_{0}(\cdot\mid\mathbf{l})$ (e.g., continuity in $a$), the grid-based approximation error in cutoff and membership is controlled by these two terms: a stochastic part ($\Delta_{n}$) and a deterministic discretization part ($\delta_{m}$). Heuristically,
$\text{HDR approximation error}\;\lesssim\;\Delta_{n}(\mathbf{l})+\delta_{m},$
uniformly in $\mathbf{l}$ under corresponding uniform conditions.
Hence, approximation improves when (i) $\hat{f}_{n}$ is accurate and (ii) the grid is sufficiently fine. In practice, one can assess adequacy by a simple grid-refinement check: recompute $\hat{f}_{n,\alpha}$, $\hat{\tau}_{n,\alpha}$, and the feasible curve on progressively finer grids (e.g., halve $\delta_{m}$) and verify numerical stability. Large changes under refinement indicate that the current grid is too coarse in regions relevant to the estimated cutoff.
Appendix D Proof of Convergence of the Feasible Curve
D.1 Setup, notation, and target
Let $O=(\mathbf{L},A,Y)\sim P_{0}$ and let $\{O_{i}\}_{i=1}^{n}$ be i.i.d. draws from $P_{0}$. Denote by $f_{0}(a\mid\mathbf{l})$ the conditional density of $A$ given $\mathbf{L}=\mathbf{l}$ and by $\mu_{0}(a,\mathbf{l})=\mathbb{E}_{0}(Y\mid A=a,\mathbf{L}=\mathbf{l})$ the outcome regression.
Fix a support level $\alpha\in(0,1)$. For each $\mathbf{l}$, let
$\mathcal{A}_{\alpha}(\mathbf{l};P_{0})=\{a\in\mathcal{A}:\ f_{0}(a\mid\mathbf{l})\geq f_{\alpha}(\mathbf{l})\}$
be the $\alpha$-HDR region defined in Appendix C. Define the feasible intervention operator
$d_{\alpha}(a,\mathbf{l})=\begin{cases}a,&a\in\mathcal{A}_{\alpha}(\mathbf{l};P_{0}),\\ \Pi_{\mathcal{A}_{\alpha}(\mathbf{l};P_{0})}(a),&a\notin\mathcal{A}_{\alpha}(\mathbf{l};P_{0}),\end{cases}$
where $\Pi_{T}(\cdot)$ is the (single-valued) projection onto a nonempty closed set $T\subset\mathbb{R}$ using a fixed deterministic measurable tie-breaking convention.
Under the identification conditions stated in the main text, the feasible curve can be written as